# Empirical Dataset for Image Transformations

### v1

@@ Abstract
Existing image editing models struggle to meet real-world demands; despite excelling in academic benchmarks, we are yet to see them adopted to solve real user needs. The datasets that power these models use artificial edits, lacking the scale and ecological validity necessary to address the true diversity of user requests. In response, we introduce RealEdit, a large-scale image editing dataset with authentic user requests and human-made edits sourced from Reddit. RealEdit contains a test set of $9.3$K examples the community can use to evaluate models on real user requests. Our results show that existing models fall short on these tasks, implying a need for realistic training data. So, we introduce $48$K training examples, with which we train our RealEdit model. Our model achieves substantial gains—outperforming competitors by up to $165$ Elo points in human judgment and $92\%$ relative improvement on the automated VIEScore metric on our test set. We deploy our model back on Reddit, testing it on new requests, and receive positive feedback. Beyond image editing, we explore RealEdit ’s potential in detecting edited images by partnering with a deepfake detection non-profit. Finetuning their model on RealEdit data improves its F1-score by $14$ percentage points, underscoring the dataset’s value for broad, impactful applications.
@@ 1 Introduction
Our model demonstrates significantly better performance than existing state-of-the-art models on RealEdit’s test set with a human preference (N=4,196) Elo score of $1184$, beating the next best model by $165$ points. We also outperform existing models using automated metrics: our model achieves $4.61$ VIEScore versus the next best score of $2.4$ (amongst other metrics). Moreover, our model still remains competitive with MagicBrush and EmuEdit [65, 48] on their test set. We further validate our model by completing new Reddit requests and receiving positive feedback.
Finally, we partner with <REDACTED>, a non-profit aimed at AI-generated content detection. By adding human-made edits from our dataset, we improve their model’s F1-score by $14$ points. Our ecologically valid experimentation highlights the dataset’s value outside of editing tasks.
@@ 3.1 Dataset creation pipeline
For the RealEdit test set, we generated captions for both input and edited images using vision-language models to support evaluation on caption-based metrics. Implementation details are provided in the Appendix.
Step 3: Data verification and final composition. After generating the dataset, we conducted a rigorous multi-stage verification process to ensure data quality. All images were screened for inappropriate content using the opennsfw2[4] network, filtering out those flagged as explicit. Additionally, RealEdit test set was manually reviewed by two annotators evaluating the following criteria: (1) appropriateness of the input image, (2) applicability of the instruction, and (3) correctness of the output image. Approximately 78% of the data points were agreed upon as high quality and included in the final test set for RealEdit.
@@ 4 RealEdit dataset analysis
We create a taxonomy of image editing tasks people have requested. This involves (1) categorizing our edit requests into operations, (2) subcategorizing requests by subject of the edit, and (3) prompting GPT-4o [39] to categorize the request based on the input image and edit instruction. To determine the operations, we modify the MagicBrush [65] set of operations for clarity. We base our possible subjects on the significant categories from the sample of 500 images. We tune our GPT-4o prompt using samples of 100 data points to ensure accuracy, then validate on a separate sample to avoid overfitting. We find that both the categories and their distribution differ greatly from prior work. Since our categorizations are fairly similar to MagicBrush [65] and Emu Edit [48] test sets, we run our taxonomy on these test sets and highlight key distributional differences in Figure 4. Full taxononomies and comparisons are listed in Appendix.
@@
Synthetic datasets contain a greater use of “add” requests (36% less than MagicBrush than). In contrast, real-life photos typically contain the intended objects within the frame, with many of RealEdit’s semantically focused tasks involving the removal of unintended elements, such as strangers in the background, shadows on faces, or cars on the street. Additionally, there are numerous cases where input images are semantically aligned with the owner’s intent, but errors in photography such as bad lighting, motion blur, or graininess. Following this, RealEdit contains more “enhance” requests (14% greater than MagicBrush and Emu Edit) compared to existing datasets. These findings indicate that real users often prioritize subtler requests, whereas synthetic datasets are dominated by larger semantic changes, such as “add.”
@@
Analysis on 500 samples reveals that around 55% of the input images feature people as the main subject. Consequently, the subjects of the requested edits are more likely to be people (13% more than Emu Edit), and less likely to be man-made objects (20% less than MagicBrush). Animals and media (characters, movie/book posters, memes, etc.) are the next most common categories, comprising about 10% of the test set each. Common media requests include restoring old photographs, participating in fandoms, making memes, or other forms of online entertainment. The fixation on media is not paralleled in other datasets (15% more than Emu Edit). These findings reveal a clear difference: Reddit users tend to prioritize personal significance by including people and entertainment by incorporating media, and synthetic datasets often fail to reflect these preferences accurately.
@@ 5 An editing model trained with RealEdit
To demonstrate the value of RealEdit, we develop an image editing model using training examples in RealEdit. Specifically, we utilize InstructPix2Pix [6] as the model backbone on which we finetune using our data. We leave exploration on using different base models as future work.
@@
Since we finetune InstructPix2Pix rather than training the model from scratch, we align our finetuning dataset with the data distribution used in InstructPix2Pix’s pretraining data to avoid substantial distributional shifts that may deteriorate model’s performance. In particular, InstructPix2Pix [6] applies CLIP-based [44] filtering to ensure the quality of image pairs. In addition, as it employs Prompt-to-Prompt [20] in generating its training data, the input-output image pairs are with high structural similarity. To align our training set, we thus follow the same CLIP-based filtering and additionally use SSIM [55] to include structurally similar images, recognizing that human edits collected in RealEdit often alter structure with techniques like drag-and-drop adjustments and symmetrical flipping. Tasks incompatible with InstructPix2Pix’s capabilities, such as resizing images, changing file types, or highly ambiguous prompts (particularly those involving humor) are thus excluded. In total, we trained on 39K examples. Aligning our training data with the InstructPix2Pix distribution allows for more competitive performance on metrics, and accounts for limitations in the InstructPix2Pix’s architecture and pretraining. For training our model, we closely follow the configuration of MagicBrush [65]. Specifically, we train our model for 51 epochs, utilizing cosine learning rate decay and incorporating a learning rate warm-up phase (details in Appendix).
@@ 6.4 Evaluations on existing test sets
We also conduct evaluations on external test sets including the test sets in GenAI Arena [23], Emu Edit [48], and MagicBrush [65]. On GenAI Arena, we report Elo ranking in Table 3 computed with real human preferences. Our model ranks second among the evaluated models. While these results were insightful, we found the examples in GenAI Arena to be less representative of real-world tasks. We include full automated evaluation results on Emu Edit and MagicBrush in Appendix, where our model performs competitively with the individual strongest models on respective test sets across varying metrics.
@@ 6.5 Improving edited image detection
Table 4: Model performance on detecting edited images in the RealEdit test set and in-the-wild images on <REDACTED>.
@@
We partnered with <REDACTED>, a platform where users can upload media to assess authenticity. Their primary fake image detection model is a fine-tuned version of Universal Fake Detect (UFD)  [38], which effectively detects model-generated deepfakes. We leverage the human-edited images in RealEdit to enhance the model’s ability to detect such edits, which has significant real-world impact.
UFD is trained on a recipe of 62K images from academic datasets [56, 25, 26, 24, 31, 51] and some proprietary data, none of which includes human edits. We trained a model from scratch using the UFD training pipeline with added RealEdit training data. We evaluated on the RealEdit test set and on a random subset of 100 reals plus 100 in-the-wild edited images from <REDACTED>. We show in Table 4 that fine-tuning on RealEdit improves F1 by 45.5 and 14 points on RealEdit and <REDACTED>’s test sets respectively. RealEdit also serves as a challenging human-edited image detection benchmark for models that are more specialized for this task compared to deepfake detection [52, 67, 12, 68].
@@ 7 Discussion
To protect user privacy, individuals can opt out of having their images in the dataset by removing the photos from Reddit. Since our dataset contains image URLs rather than image files, images deleted from the web are automatically removed. Additionally, we provide a form where individuals can request their data to be removed from the dataset. Although this evolving dataset may introduce challenges for quantitative validation, ensuring user privacy remains our top priority.
Our work has positive social impacts, such as reducing the need for professional editing software and skills, and enabling higher-quality restorations of family photographs. However, we recognize the risks of malicious exploitation and strongly oppose any harmful, offensive, or derogatory use of our model or data. We plan to further pursue the development of fake image detection tools.
Conclusion.
We propose RealEdit, a dataset of 57K input - instruction - outputs data points where all instructions and edits are performed by humans. We analyze the distribution of real-world editing requests and fine-tune InstructPix2Pix [6] to create a SOTA image editing model on these tasks. Lastly we explore RealEdit ’s potential in facilitating deepfake detection.
@@ Appendix B Data processing
Appendix C Discussion
C.1 Limitations and future work
@@ C.2 Social impacts
Accessible image editing models that are capable of handling real world tasks are extremely useful in democratizing the documentation of people’s lives. For example, some requests in RealEdit involve restoring old photographs, many of which were paid. The RealEdit model can help more users to document meaningful family histories, even if they cannot afford to pay for edits. We have demonstrated the efficacy of our model on making real world edits by uploading our model’s generations to Reddit. Additionally, our exploration of the contribution of RealEdit in deepfake image detection has shown that RealEdit increases the ability of <REDACTED>’s ability to detect fake images, which is extremely useful in a world where images are routinely edited to cause scandals or spread misinformation.
@@
C.3 Ethics
Some other editing datasets [65] do not use human faces in order to evade biases as well as privacy concerns. However, in RealEdit, we determine that since over half of edit requests contain images focused on people, we must train on human data in order to be successful in completing real world editing tasks. To mitigate privacy concerns, we use the URL in place of the actual input image so that if the original poster (OP) deletes their post, it will be removed from our dataset. We also include a form for users to request their data to be removed. In the case of mitigating biases, we hope in future work to study the effects of using Reddit data on task completion for a wide array of demographic groups, as well as techniques or supplementary data sources to boost performance on underrepresented groups. This is a known problem in the field, and we are compelled by user preferences to include human data. Given this, although we appreciate the importance of mitigating demographic biases, this is outside the scope of a single paper.
Appendix D Modeling ablations
D.1 Implementation details
@@ D.4 Processing instructions
Appendix E Inference time results
E.1 Hyperparameters
@@ E.2 Instruction rewriting
E.3 Quantitative evaluation on external test sets
@@ E.4 Elo scores
Appendix F Reddit experiment
@@
Appendix G Edited image detection
@@
The baseline classifier undergoes a multi-stage training process: initially on academic datasets and subsequently fine-tuned on <REDACTED>’s proprietary data. In total, the baseline model is trained on 65K images with a near equal 50/50 split between real and generated images. To assess the value of RealEdit data for fake image detection, we train a second version of UFD by combining the original data with RealEdit data. Specifically, we include only photographs, excluding non-photographic images such as digital artworks, screenshots, cartoons, and infographics, filtered using GPT-4o. This single-stage training incorporates an additional 37K original and 37K edited images, resulting in a total of 139K images.
In the first stage of training, the <REDACTED> model took over 24 hours to train on an A10G GPU with 20GB of RAM and the remaining three stages took 4 hours. Our optimized model took 1.5 hours to train on a L40S GPU with 40GB of RAM.
Table 11: Breakdown of fake image sources in the training recipe of the <REDACTED> model used as our baseline.
@@
Fakes uploaded to <REDACTED> 2K
@@
Reals uploaded to <REDACTED> 0.7K
<REDACTED>’s in-the-wild test set
<REDACTED>’s in-the-wild test set includes images uploaded between 8/16/2024 and 11/10/2024. We randomly sample 100 real images and then sample 100 fake images selected from those tagged as ”likely photoshopped” by professional sleuths in <REDACTED>’s media database, ensuring the evaluation focuses on human-edited images rather than exclusively AI-generated edits. Tool usage data was available for some images, revealing that approximately 80% of the fake images were human edits created with Photoshop, while the remaining 20% involved human edits combined with AI tools such as Dream Studio AI, Insightface AI, and Remaker AI.
@@
Appendix H Additional results

### v2

@@ Abstract
Existing image editing models struggle to meet real-world demands; despite excelling in academic benchmarks, we are yet to see them adopted to solve real user needs. The datasets that power these models use artificial edits, lacking the scale and ecological validity necessary to address the true diversity of user requests. In response, we introduce RealEdit, a large-scale image editing dataset with authentic user requests and human-made edits sourced from Reddit. RealEdit contains a test set of $9.3$K examples the community can use to evaluate models on real user requests. Our results show that existing models fall short on these tasks, implying a need for realistic training data. So, we introduce $48$K training examples, with which we train our RealEdit model. Our model achieves substantial gains—outperforming competitors by up to $165$ Elo points in human judgment and $92\%$ relative improvement on the automated VIEScore metric on our test set. We deploy our model back on Reddit, testing it on new requests, and receive positive feedback. Beyond image editing, we explore RealEdit ’s potential in detecting edited images by partnering with a deepfake detection non-profit. Finetuning their model on RealEdit data improves its F1-score by $14$ percentage points, underscoring the dataset’s value for broad, impactful applications.  [^1] Project page: https://peter-sushko.github.io/RealEdit
@@ 1 Introduction
Our model demonstrates significantly better performance than existing state-of-the-art models on RealEdit’s test set with a human preference (N=4,196) Elo score of $1184$, beating the next best model by $165$ points. We also outperform existing models using automated metrics: our model achieves $4.61$ VIEScore versus the next best score of $2.4$ (amongst other metrics). Moreover, our model still remains competitive with MagicBrush and EmuEdit [67, 50] on their test sets. We further validate our model by completing new Reddit requests and receiving positive feedback.
Finally, we partner with TrueMedia.org, a non-profit aimed at AI-generated content detection. By adding human-made edits from our dataset, we improve their model’s F1-score by $14$ points. Our ecologically valid experimentation highlights the dataset’s value outside of editing tasks.
@@ 3.1 Dataset creation pipeline
For the RealEdit test set, we generated captions for both input and edited images using vision-language models to support evaluation on caption-based metrics. Implementation details are provided in Appendix B.
Step 3: Data verification and final composition. After generating the dataset, we conducted a rigorous multi-stage verification process to ensure data quality. All images were screened for inappropriate content using the opennsfw2 [4] network, filtering out those flagged as explicit. Additionally, the RealEdit test set was manually reviewed by two annotators evaluating the following criteria: (1) appropriateness of the input image, (2) applicability of the instruction, and (3) correctness of the output image. We employed qualitative coding, establishing four distinct codes, with Cohen’s Kappa yielding 0.61, indicating substantial agreement. Approximately 78% of the data points were agreed upon as high quality and included in the final test set for RealEdit, mitigating individual biases.
@@ 4 RealEdit dataset analysis
We create a taxonomy of image editing tasks people have requested. This involves (1) categorizing our edit requests into operations, (2) subcategorizing requests by subject of the edit, and (3) prompting GPT-4o [41] to categorize the request based on the input image and edit instruction. To determine the operations, we modify the MagicBrush [67] set of operations for clarity. We base our possible subjects on the significant categories from the sample of 500 images. We tune our GPT-4o prompt using samples of 100 data points to ensure accuracy, then validate on a separate sample to avoid overfitting. We find that both the categories and their distribution differ greatly from prior work. Since our categorizations are fairly similar to MagicBrush [67] and Emu Edit [50] test sets, we run our taxonomy on these test sets and highlight key distributional differences in Figure 4. Full taxonomies and comparisons are listed in Appendix A.
@@
Synthetic datasets contain a greater use of “add” requests (36% less than MagicBrush). In contrast, real-life photos typically contain the intended objects within the frame, with many of RealEdit’s semantically focused tasks involving the removal of unintended elements, such as strangers in the background, shadows on faces, or cars on the street. Additionally, there are numerous cases where input images are semantically aligned with the owner’s intent, but contain errors in photography such as bad lighting, motion blur, or graininess. Following this, RealEdit contains more “enhance” requests (14% greater than MagicBrush and Emu Edit) compared to existing datasets. These findings indicate that real users often prioritize subtler requests, whereas synthetic datasets are dominated by larger semantic changes, such as “add.”
@@
Analysis on 500 samples reveals that around 55% of the input images feature people as the main subject. Consequently, the subjects of the requested edits are more likely to be people (13% more than Emu Edit), and less likely to be man-made objects (20% less than MagicBrush). Animals and media (characters, movie/book posters, memes, etc.) are the next most common categories, comprising about 10% of the test set each. Common media requests include restoring old photographs, participating in fandoms, making memes, or other forms of online entertainment. The fixation on media is not paralleled in other datasets (15% more than Emu Edit). These findings reveal a clear difference: Reddit users tend to prioritize personal significance by including people and entertainment by incorporating media. Synthetic datasets often fail to reflect these preferences accurately.
@@ 5 An editing model trained with RealEdit
To demonstrate the value of RealEdit, we develop an image editing model using training examples in RealEdit. Specifically, we utilize InstructPix2Pix [6] as the model backbone on which we finetune using our data. We leave exploration of different base models for future work.
@@
Since we finetune InstructPix2Pix rather than training the model from scratch, we align our finetuning dataset with the data distribution used in InstructPix2Pix’s pretraining data to avoid substantial distributional shifts that may deteriorate model’s performance. In particular, InstructPix2Pix [6] applies CLIP-based [46] filtering to ensure the quality of image pairs. In addition, as it employs Prompt-to-Prompt [21] in generating its training data, the input-output image pairs are with high structural similarity. To align our training set, we thus follow the same CLIP-based filtering and additionally use SSIM [57] to include structurally similar images, recognizing that human edits collected in RealEdit often alter structure with techniques like drag-and-drop adjustments and symmetrical flipping. Tasks incompatible with InstructPix2Pix’s capabilities, such as resizing images, changing file types, or highly ambiguous prompts (particularly those involving humor) are thus excluded. In total, we trained on 39K examples. Aligning our training data with the InstructPix2Pix distribution allows for more competitive performance on metrics, and accounts for limitations in the InstructPix2Pix’s architecture and pretraining. For training our model, we closely follow the configuration of MagicBrush [67]. Specifically, we train our model for 51 epochs, utilizing cosine learning rate decay and incorporating a learning rate warm-up phase (details in Appendix E.
@@ 6.4 Evaluations on existing test sets
We also conduct evaluations on external test sets including the test sets in GenAI Arena [25], Emu Edit [50], and MagicBrush [67]. On GenAI Arena, we report Elo ranking in Table 3 computed with real human preferences. Our model ranks second among the evaluated models. While these results were insightful, we found the examples in GenAI Arena to be less representative of real-world tasks. Additionally, since GenAI Arena’s data comes from ImagenHub, which includes MagicBrush data, the evaluation is inherently biased toward MagicBrush. We include full automated evaluation results on Emu Edit and MagicBrush in Appendix F, where our model performs competitively with the individual strongest models on respective test sets across varying metrics.
@@ 6.5 Improving edited image detection
Table 4: Model performance on detecting edited images in the RealEdit test set and in-the-wild images on TrueMedia.org.
@@
We partnered with TrueMedia.org, a platform where users can upload media to assess authenticity. Their primary fake image detection model is a fine-tuned version of Universal Fake Detect (UFD)  [40], which effectively detects model-generated deepfakes. We leverage the human-edited images in RealEdit to enhance the model’s ability to detect such edits, which has significant real-world impact.
UFD is trained on a recipe of 62K images from academic datasets [58, 27, 28, 26, 33, 53] and some proprietary data, none of which includes human edits. We trained a model from scratch using the UFD training pipeline with added RealEdit training data. We evaluated on the RealEdit test set and on a random subset of 100 reals plus 100 in-the-wild edited images from TrueMedia.org. We show in Table 4 that fine-tuning on RealEdit improves F1 by 45.5 and 14 points on RealEdit and TrueMedia.org’s test sets respectively. RealEdit also serves as a challenging human-edited image detection benchmark for models that are more specialized for this task compared to deepfake detection [54, 69, 13, 70].
@@ 7 Discussion
To protect user privacy, individuals can opt out of having their images in the dataset by removing the photos from Reddit. Since our dataset contains image URLs rather than image files, images deleted from the web are automatically removed. Additionally, we provide a form where individuals can request their data to be removed from the dataset.
Our work has positive social impacts, such as reducing the need for professional editing software and skills, and enabling higher-quality restorations of family photographs. However, we recognize the risks of malicious exploitation and strongly oppose any harmful, offensive, or derogatory use of our model or data.
Acknowledgments
We thank Galen Weld, Matthew Wallingford, Vivek Ramanujan, and Benlin Liu for their time, guidance, and mentorship. We additionally thank the UW RAIVN lab for support throughout the project.
@@ A.1 Full taxonomy
Operation %
Remove 31.9
Enhance 14.5
Add 12.5
Change Color 11.5
Restore 9.7
Replace 8.4
Open Ended 7.9
Formatting 3.4
Other 0.1
Table 5: Distribution of edit operations in the test set.
Subject %
Person 15.2
Media 14.9
Background 11.0
Body Part 9.3
Text 8.9
Object 8.3
Apparel 6.5
Formatting 3.4
Animal 3.3
Logo 2.5
Vehicle 2.4
Nature 2.2
Other 12.1
Table 6: Distribution of edit subjects in the test set.
@@ Appendix B Data processing
Appendix C Additional baselines
After the submission of this paper, we were made aware of additional editing models with strong performance. In Table 8, we compare RealEdit to Edit Friendly [24] and TurboEdit [11].
Table 8: RealEdit outperforms Edit Friendly and TurboEdit on real-world edit requests.
Model VSC $\uparrow$ VPQ $\uparrow$ VO $\uparrow$ L1 $\downarrow$ L2 $\downarrow$ CLIPI $\uparrow$ DINOI $\uparrow$ CLIPT $\uparrow$
Turbo-Edit 2.73 5.19 2.80 0.147 0.066 0.782 0.739 0.261
Edit Friendly 3.00 5.68 2.93 0.199 0.091 0.762 0.736 0.261
RealEdit 3.68 4.01 4.61 0.143 0.066 0.840 0.792 0.261
@@ Appendix C Additional baselines
Appendix D Discussion
D.1 Limitations and future work
@@ D.2 Social impacts
Accessible image editing models that are capable of handling real world tasks are extremely useful in democratizing the documentation of people’s lives. For example, some requests in RealEdit involve restoring old photographs, many of which were paid. The RealEdit model can help more users to document meaningful family histories, even if they cannot afford to pay for edits. We have demonstrated the efficacy of our model on making real world edits by uploading our model’s generations to Reddit. Additionally, our exploration of the contribution of RealEdit in deepfake image detection has shown that RealEdit increases the ability of TrueMedia.org’s ability to detect fake images, which is extremely useful in a world where images are routinely edited to cause scandals or spread misinformation.
@@
D.3 Ethics
Some other editing datasets [67] do not use human faces in order to evade biases as well as privacy concerns. However, in RealEdit, we determine that since over half of edit requests contain images focused on people, we must train on human data in order to be successful in completing real world editing tasks. To mitigate privacy concerns, we use the URL in place of the actual input image so that if the original poster (OP) deletes their post, it will be removed from our dataset. We also include a form for users to request their data to be removed. This follows the standards set by RedCaps  [10]. In the case of mitigating biases, we hope in future work to study the effects of using Reddit data on task completion for a wide array of demographic groups, as well as techniques or supplementary data sources to boost performance on underrepresented groups. This is a known problem in the field, and we are compelled by user preferences to include human data. Given this, although we appreciate the importance of mitigating demographic biases, this is outside the scope of a single paper.
Appendix E Modeling ablations
E.1 Implementation details
@@ E.4 Processing instructions
Appendix F Inference time results
F.1 Hyperparameters
@@ F.2 Instruction rewriting
F.3 Quantitative evaluation on external test sets
@@ F.4 Elo scores
Appendix G Reddit experiment
@@
Appendix H Edited image detection
@@
The baseline classifier undergoes a multi-stage training process: initially on academic datasets and subsequently fine-tuned on TrueMedia.org’s proprietary data. In total, the baseline model is trained on 65K images with a near equal 50/50 split between real and generated images. To assess the value of RealEdit data for fake image detection, we train a second version of UFD by combining the original data with RealEdit data. Specifically, we include only photographs, excluding non-photographic images such as digital artworks, screenshots, cartoons, and infographics, filtered using GPT-4o. This single-stage training incorporates an additional 37K original and 37K edited images, resulting in a total of 139K images.
In the first stage of training, the TrueMedia.org model took over 24 hours to train on an A10G GPU with 20GB of RAM and the remaining three stages took 4 hours. Our optimized model took 1.5 hours to train on a L40S GPU with 40GB of RAM.
Table 14: Breakdown of fake image sources in the training recipe of the TrueMedia.org model used as our baseline.
@@
Fakes uploaded to TrueMedia.org 2K
@@
Reals uploaded to TrueMedia.org 0.7K
TrueMedia.org’s in-the-wild test set
TrueMedia.org’s in-the-wild test set includes images uploaded between 8/16/2024 and 11/10/2024. We randomly sample 100 real images and then sample 100 fake images selected from those tagged as ”likely photoshopped” by professional sleuths in TrueMedia.org’s media database, ensuring the evaluation focuses on human-edited images rather than exclusively AI-generated edits. Tool usage data was available for some images, revealing that approximately 80% of the fake images were human edits created with Photoshop, while the remaining 20% involved human edits combined with AI tools such as Dream Studio AI, Insightface AI, and Remaker AI.
@@
Appendix I Additional results